40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4802     Entry time: Thu Jun 9 20:10:38 2011
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: ETMY whitening filter : all off 

I checked the state of the whitening filters for the ETMY shadow sensors.

Result : They've been OFF  (i.e. flat response).

 

(measurement and setup)

 I measured the transfer functions of the whitening board (D000210) by looking at the signal before and after the whitening stage.

 The whitening board handles five signals; UL, UR, LR, LL and SD, and there are five single-pin lemo outputs for each signal on the front panel.

A good thing on those lemo monitors is that their signals are monitored before the whitening stages.

Rana suggested me to use these signals for the denominator of the transfer functions and consider the sensor signals as excitation signals.

So I plugged those signals into extra ADC channels via an AA-board and measured the transfer functions.

In the measurement the coherence above 4 Hz was quite small while the suspension was freely swinging.

Therefore I had to excite the ETMY suspension by putting random noise in a frequency band from 5 Hz to 35 Hz to obtain better coherence.

 

(results)

 The response is flat over frequency range from ~ 0.2 Hz to ~40 Hz, see the plot below. 

According to the spectrum of each signal the measurements above 10 Hz are just disturbed by the ADC noise.

If the whitening filters are ON, a pole and zero are expected to appear at 30 Hz and 3 Hz respectively according to the schematic, but no such features.

ETMY_WF2.png

 

ELOG V3.1.3-